a body travels half of its total path in the last second of its free fall fron rest. The duration of fall is nearly

Grade:9

2 Answers

Sher Mohammad IIT Delhi
9 years ago
velocity of the body when it travels half of the path v=sqrt(g*h)
Now in 1 second it travels rest half path so s=u*t+.5*g*t^2
0.5h=sqrt(gh)+0.5*g
=>h^2-6gh+g^2=0
=>h=.5*g*(6+4sqrt(2))

now t=sqrt(2*h/g)
t=sqrt(6+4sqrt(2))

Sher mohammad
askiitians faculty
Arun Kumar IIT Delhi
9 years ago



Thanks & Regards
Arun Kumar
IIT Delhi
Askiitians Faculty